OpenStudy (anonymous):

how do you rewrite using only positive exponents -3m^-2n^-4

OpenStudy (shane_b):

So it's this? \[\large -3m^{-2}n^{-4}\]

OpenStudy (anonymous):

oops yeah im wrong. wrong typo. -3/(m^2n^4)

OpenStudy (shane_b):

If it's this:\[\large -3m^{-2}n^{-4}\]then you can simply convert it to a fraction with positive denominators:\[\large \frac{-3}{m^{2}n^{4}}\]

OpenStudy (shane_b):

(that should read: positive exponents in the denominators)

jhonyy9 (jhonyy9):

do you understand it ? because x^(-1) = 1/x ok ?